Are the 3 below written correctly in Algebraic symbols?
Three more than the sum of x and y = x+y=3
c subtracted from the sum of x and y = x+y=c
If one addend of x is c, what is the other addend? b
Please verify if these 3 word problems above have been written in correct Algebraic symbols?
Thank you


Answer by Alan3354(69443)

Three more than the sum of x and y = x+y=3
The sum of x & y is x+y
3 more than that is x+y + 3
--------------------------------------
c subtracted from the sum of x and y = x+y=c
x+y - c
=========================
If one addend of x is c, what is the other addend? b
x-c
=================================
Please verify if these 3 word problems above have been written in correct Algebraic symbols?
Thank you